OpenStudy (anonymous):

if the point P(-4/5,y) is on the unit circle and P is in the 3rd quadrant, what is y equal to?

OpenStudy (anonymous):

you know \((-\frac{4}{5})^2+y^2=1\) so you can solve for \(y\)

OpenStudy (anonymous):

thanks :) i'll solve for it

OpenStudy (anonymous):

or you can remember the 3 - 4 - 5 right triangle

OpenStudy (anonymous):

i got y = (3/5), (-3/5)

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):



OpenStudy (anonymous):

yes it is right, but there is only one answer, not two

OpenStudy (anonymous):

or is y only equal to one of those values?

OpenStudy (anonymous):

right, i have to check which one i forgot

OpenStudy (anonymous):

because you are in quadrant 3 you the second coordinate is below the \(x\) axis and so it is negative

OpenStudy (anonymous):

so y = (3/5)

OpenStudy (anonymous):

oh sorry i didn't know it had to be negative

OpenStudy (anonymous):

so in that case y would equal (-3/5)?

OpenStudy (anonymous):

yes it is negative, but this should not be a mystery if you know what quadrant you are in 

OpenStudy (anonymous):

thanks, i forgot about the quadrant. thanks for the diagram

OpenStudy (anonymous):

left and down, means both coordinates of that point are negative

OpenStudy (anonymous):

thanks! i understand it much better now. can you please help me with another problem similar to this one? i'd really appreciate it, i'll post it now

OpenStudy (anonymous):

yw (clear right?)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

yep, very clear :) thanks! i just posted the other one as well